### v1

# Hybrid Quantum-Classical Recurrent Neural Networks

Abstract
We present a hybrid quantum-classical recurrent neural network (QRNN) architecture in which the entire recurrent core is realized as a parametrized quantum circuit (PQC) controlled by a classical feedforward network. The hidden state is the quantum state of an $n$-qubit PQC, residing in an exponentially large Hilbert space $\mathbb{C}^{2^{n}}$. The PQC is unitary by construction, making the hidden-state evolution norm-preserving without external constraints. At each timestep, mid-circuit readouts are combined with the input embedding and processed by the feedforward network, which provides explicit classical nonlinearity. The outputs parametrize the PQC, which updates the hidden state via unitary dynamics. The QRNN is compact and physically consistent, and it unifies (i) unitary recurrence as a high-capacity memory, (ii) partial observation via mid-circuit measurements, and (iii) nonlinear classical control for input-conditioned parametrization. We evaluate the model in simulation with up to 14 qubits on sentiment analysis, MNIST, permuted MNIST, copying memory, and language modeling, adopting projective measurements as a limiting case to obtain mid-circuit readouts while maintaining a coherent recurrent quantum memory. We further devise a soft attention mechanism over the mid-circuit readouts in a sequence-to-sequence model and show its effectiveness for machine translation. To our knowledge, this is the first model (RNN or otherwise) grounded in quantum operations to achieve competitive performance against strong classical baselines across a broad class of sequence-learning tasks.
1 Introduction
Recurrent neural networks (RNNs) process sequence data by maintaining a hidden state that is updated at each timestep, which can create a bottleneck for memory and representational capacity. While vanilla RNNs have been empirically shown to retain roughly one real value of information per hidden unit, with the effective task-specific capacity linearly bounded by the number of model parameters [Collins et al., 2017], similar limitations extend to gated architectures such as LSTMs and GRUs [Hochreiter and Schmidhuber, 1997, Cho et al., 2014], despite their use of gating and explicit memory cells [Collins et al., 2017]. This means that more complex sequences may exceed what the hidden state can encode, forcing the model to compress or forget.
The introduction of the Transformer [Vaswani et al., 2017] appeared to obviate explicit recurrence by bypassing the hidden-state bottleneck. However, recent work shows that recurrent inductive bias remains highly competitive and provides representational advantages not matched by Transformers [Gu and Dao, 2023, Orvieto et al., 2023, Bhattamishra et al., 2024, Beck et al., 2024].
Another challenge in training RNNs is the vanishing and exploding gradient problem [Bengio et al., 1994, Hochreiter and Schmidhuber, 1997], which arises from repeated multiplication through the recurrent Jacobian. Among various strategies to address this [Mikolov, 2012, Pascanu et al., 2013, Le et al., 2015], unitary and orthogonal RNNs [Arjovsky et al., 2016, Jing et al., 2019, Helfrich et al., 2018, Kiani et al., 2022] constrain the recurrent weights to be norm-preserving, allowing gradients to remain stable across timesteps. These models perform well on synthetic tasks, but their results on broader benchmarks vary.
With the advancement of quantum computing [Arute et al., 2019, Acharya et al., 2024, Reichardt et al., 2024, DeCross et al., 2025], parametrized quantum circuits (PQCs), which are a core component of variational hybrid quantum-classical models, have concurrently emerged as an alternative mechanism for function approximation [Benedetti et al., 2019, Du et al., 2019, Bondesan and Welling, 2020, Pérez-Salinas et al., 2021, Schuld et al., 2021, Yu et al., 2024b]. PQCs implement unitary transformations by construction, which naturally preserve norms (§3.1). Acting on $n$ qubits, they enable expressive transformations over quantum states in an exponentially large Hilbert space $\mathbb{C}^{2^{n}}$. Although such spaces are classically intractable beyond moderate $n$, they can be manipulated with only $n$ qubits on quantum hardware.
$\framebox{$\mathbf{RX}$}\raisebox{-6.02773pt}{\hskip 1.72218pt$\scriptstyle i$}\;=\;{\begin{bmatrix}\cos\!\left(\frac{\mathstrut\theta_{i}}{\mathstrut 2}\right)&-i\sin\!\left(\frac{\mathstrut\theta_{i}}{\mathstrut 2}\right)\\ -i\sin\!\left(\frac{\mathstrut\theta_{i}}{\mathstrut 2}\right)&\cos\!\left(\frac{\mathstrut\theta_{i}}{\mathstrut 2}\right)\end{bmatrix}}\text{,}\quad\framebox{$\mathbf{RY}$}\raisebox{-6.02773pt}{\hskip 1.72218pt$\scriptstyle i$}\;=\;{\begin{bmatrix}\cos\!\left(\frac{\mathstrut\theta_{i}}{\mathstrut 2}\right)&-\sin\!\left(\frac{\mathstrut\theta_{i}}{\mathstrut 2}\right)\\ \sin\!\left(\frac{\mathstrut\theta_{i}}{\mathstrut 2}\right)&\cos\!\left(\frac{\mathstrut\theta_{i}}{\mathstrut 2}\right)\end{bmatrix}}$.
Figure 1: Hybrid QRNN. (a) Recurrent core PQC with $n=4$ qubits (illustrative) and 16 parametrized gates, acting on a quantum state in the Hilbert space $\mathbb{C}^{2^{n}}$; each horizontal line corresponds to one qubit. (b) $\mathbf{RX}$ and $\mathbf{RY}$ gates.22footnotemark: 2 Each gate in the PQC is parametrized by a rotation angle $\theta_{i}$, where $1\leq i\leq 16$. (c) QRNN unrolled for a sequence of length $k$. The feedforward network $\mathcal{F}$ takes as input the concatenation of the mid-circuit readout vector from the previous timestep and the current input $(\mathbf{z}_{t-1}\mathpunct{:}\mathbf{x}_{t})$, and outputs $\bm{\theta}_{t}\in\mathbb{R}^{16}$ containing the 16 rotation angles that parametrize the PQC shown in (a) as $U(\bm{\theta}_{t})$, which acts on the quantum state propagated from the previous step. Refer to caption denotes qubit measurements.
In this work, we present a hybrid quantum–classical recurrent neural network (QRNN) architecture grounded in quantum operations, in which the entire recurrent core is realized as a PQC. The hidden state is the quantum state of the PQC, residing in an exponentially large Hilbert space. A classical feedforward network parametrizes and steers the quantum computation, introducing nonlinearity through mid-circuit readouts, which are realized in simulation as a limiting case via projective measurements. This avoids emulating nonlinearity through linear quantum dynamics, leaving the PQC strictly for coherent unitary evolution.
Fig. 2 illustrates both the PQC (with four qubits shown for illustration) and the unrolled QRNN:
  • At each timestep $t$, the input is mapped to a classical embedding $\mathbf{x}_{t}$ via a learnable embedding layer.
  • A classical feedforward network $\mathcal{F}$ takes as input the concatenation of the readout vector $\mathbf{z}_{t-1}$ (outputs from all measurements at timestep t1t-1) and the current input $\mathbf{x}_{t}$. It outputs the PQC parameters $\bm{\theta}_{t}$ which configure the PQC with a fixed gate layout (Fig. 1(a)), denoted $U(\bm{\theta}_{t})$, applied at timestep $t$ (Fig. 1(c)).
  • The PQC applies the parametrized unitary gates to evolve the quantum state, yielding the updated state. Residing in an exponentially large Hilbert space, this state persists across timesteps and provides the model’s core recurrent memory.
  • The mid-circuit readout $\mathbf{z}_{t}$ (or the final readout at the end of the sequence) is a classical feature vector obtained from the quantum state via measurements and is used: (i) as recurrent feedback $\mathbf{z}_{t-1}$ at timestep $t$, and (ii) as the input to task-specific classical layers.
We develop the models on GPUs, allowing us to simulate and train quantum recurrence via classical backpropagation, with the expectation that such models will become classically unsimulatable as the number of qubits increases. To our knowledge, this is the first model grounded in quantum operations (RNN or otherwise) demonstrated in classical simulation with up to 14 qubits across six realistic sequence-modeling tasks, achieving competitive performance with LSTM and the scaled Cayley orthogonal scoRNN designed for norm preservation [Helfrich et al., 2018]. Experiments also show that classical nonlinear control and feedback are effective, with the nonlinear variants outperforming their linear counterparts, and that the unitary quantum recurrent core maintains more stable gradients than LSTMs (§4.6).
The QRNN is motivated in part by the memory and gradient problems of RNNs, but its main aim is to explore a hybrid quantum–classical recurrent model in an idealized proof of principle that allows us to study its computational behavior under best-case conditions across a broad class of sequence learning tasks. The PQC [Sim et al., 2019] uses only one- and two-qubit gates without nonstandard operations, and the overall architecture provides a hardware-aware base case and a plausible path toward future hardware implementations.
Another way to view the QRNN is via fast and slow weights in RNNs, which function as different types of memory across multiple timescales [Schmidhuber, 1992, Ba et al., 2016]. The PQC parameters serve as the short-term memory, analogous to the hidden activities of classical RNNs, and are controlled and reconfigured at each timestep by a classical feedforward network whose slow weights encode the long-term memory. The quantum state, updated via unitary transformations, evolves on a faster timescale than the slow weights, persists across timesteps, and acts as a third, higher-capacity memory in the Hilbert space, retaining information that influences subsequent computation [Hinton and Plaut, 1987, Schmidhuber, 1993].
2 Related Work
Bausch [Bausch, 2020] developed a QRNN with a persistent quantum memory based on quantum neurons [Cao et al., 2017]. However, nonlinearities are emulated within PQCs using probabilistic repeat-until-success circuit execution and postselection steps. The inherently linear nature of quantum computation constrains such emulation, and the available forms of nonlinearity remain limited [Yan et al., 2020, Moreira et al., 2023, Zi et al., 2024].
The so-called QLSTMs embed PQCs into the gating mechanisms of classical LSTMs [Chen et al., 2020, Yu et al., 2024a, Ubale et al., 2025], replacing dense layers in the LSTM gates with PQCs. However, all memory and recurrence remain entirely classical, governed by standard hidden and cell state updates. These architectures are best viewed as classical LSTMs augmented with auxiliary PQCs, rather than quantum recurrent models.
Li et al. [2023] and Siemaszko et al. [2023] also model recurrences with PQCs, while supporting per-timestep readouts, but they rely entirely on linear quantum dynamics of the PQC without explicit nonlinearities or classical control.
Experiments with the existing models have focused on domain-specific tasks such as fraud detection [Ubale et al., 2025], low-resource text classification [Yu et al., 2024a], or scaled-down MNIST [Bausch, 2020, Siemaszko et al., 2023]. We instead present the first QRNN to demonstrate competitive performance across six full-scale sequence modeling tasks.
3 Model
3.1 PQC
Unitary evolution.
[…]
Let $U$ denote the composition (product) of a collection of unitary gates, hence $U^{\dagger}U=I$. For any state $\lvert\psi\rangle$,
$\lVert U\lvert\psi\rangle\rVert^{2}=\langle\psi\rvert U^{\dagger}U\lvert\psi\rangle=\langle\psi\rvert I\lvert\psi\rangle=\langle\psi|\psi\rangle=\lVert\psi\rVert^{2},$
[…]
Parametrized unitary gates.
In a PQC, gates can be either fixed or parametrized. Fixed gates implement structural operations and remain constant throughout training, [^6] For example, the $\mathbf{CNOT}$ gate flips the target qubit if the control is in the $|1\rangle$ state. while the latter contain learnable parameters, which function like trainable weight matrices analogous to neural-network “layers”. The PQC in Fig. 1(a) consists of entirely parametrized gates.
Measurements.
To probe the quantum state of a PQC, we can perform measurements to obtain real-valued readouts. These readouts provide partial observations of the state, and any required number of measurements, on any of the qubits, can be combined for downstream tasks. For instance, a measurement through the Pauli-$\mathbf{Z}$ observable with the unitary
$\mathbf{Z}=\begin{pmatrix}1&0\\ 0&-1\end{pmatrix}$
assigns scalar values (e.g., +1+1 for $|0\rangle$ and 1-1 for $|1\rangle$) in the computational basis in the single-qubit case. For a general quantum state however, the outcome is probabilistic: it yields +1+1 with probability $|\alpha|^{2}$ and 1-1 with probability $|\beta|^{2}$. The expectation value of this measurement is given by $|\alpha|^{2}-|\beta|^{2}$, which can be used as a real-valued readout in hybrid quantum-classical models.
Although the readouts obtained via measurement are a nonlinear function of the gate parameters, particularly those used in parametrized rotation gates such as $\mathbf{RX}$, the resulting nonlinearity is generally weak (§4).
3.2 Hybrid Model
RNNs parameterize a conditional distribution with a function that depends on a hidden state $\mathbf{h}_{t-1}$, which compacts past inputs $(\mathbf{x}_{1},\dots,\mathbf{x}_{t-1})$ into a fixed-dimensional representation:
$p(\mathbf{x}_{t}\mid\mathbf{x}_{1},\dots,\mathbf{x}_{t-1})\;\approx\;p(\mathbf{x}_{t}\mid\mathbf{h}_{t-1}).$
At each timestep $t$, the hidden state $\mathbf{h}_{t}$ is updated based on the previous hidden state $\mathbf{h}_{t-1}$ and the current input $\mathbf{x}_{t}$:
$\mathbf{h}_{t}\;=\;f\bigl(\mathbf{h}_{t-1},\mathbf{x}_{t};\bm{\Theta}\bigr),$
where $f$ is a transformation (e.g., a basic RNN or LSTM cell) parametrized by $\bm{\Theta}$. In the hybrid model (Fig. 1(c)), we replace the hidden state with a quantum state represented by the PQC in Fig. 1(a), which is controlled by a classical feedforward network and evolved by applying the unitary gates.
Let $\mathbf{x}_{t}$ be the input embedding at timestep $t$, and let $\mathbf{z}_{t-1}$ be the measurement-based readout from the previous timestep. In the most generic form of the hybrid model, [^7] We may add extra transformations to the measurement outcomes before classifications or feeding them to the next step; see §4. the two are concatenated into a single vector $\mathbf{u}_{t}=(\mathbf{z}_{t-1}\mathpunct{:}\mathbf{x}_{t})$ and passed through a classical feedforward network $\mathcal{F}$ with one hidden layer and a nonlinearity.
The first transformation in $\mathcal{F}$ maps the input $\mathbf{u}_{t}$ to a hidden representation $\mathbf{v}_{t}$:
$\mathbf{v}_{t}=\phi(\mathbf{W}_{1}\mathbf{u}_{t}+\mathbf{b}_{1}),$ (1)
where $\phi$ is a nonlinear activation function. The second transformation maps $\mathbf{v}_{t}$ to
$\bm{\theta}_{t}=\mathbf{W}_{2}\mathbf{v}_{t}+\mathbf{b}_{2},$ (2)
where $\bm{\theta}_{t}\in\mathbb{R}^{d}$ represents the parameters that control the PQC’s unitary operations at timestep $t$. Each element of $\bm{\theta}_{t}$ denoted $\theta_{i}$ is mapped to a rotation angle in a parametrized quantum gate within the PQC (e.g., $1\leq i\leq d$ and $d=16$ in Fig. 1(a)).
The PQC itself is defined by a unitary operator $U(\bm{\theta}_{t})$, parametrized by $\bm{\theta}_{t}$. [^8] We use $U(\bm{\theta}_{t})$ to denote all unitary operations composed of multiple parametrized gates, each acting on one or more qubits with parameters drawn from $\bm{\theta}_{t}$. Applying the gates in $U(\bm{\theta}_{t})$ to the quantum state $\mathbf{h}_{t-1}=|\psi_{t-1}\rangle$ yields the updated state $\mathbf{h}_{t}=U(\bm{\theta}_{t})\,|\psi_{t-1}\rangle$. The updated state is then measured to obtain a classical readout vector
$\mathbf{z}_{t}=\textbf{Measure}(\mathbf{h}_{t}),$ (3)
which serves as a proxy for the quantum state and is combined with the next input $\mathbf{x}_{t+1}$ to evolve the recurrence. Mid-circuit measurements are simulated without collapsing the quantum state via projective measurements, thereby preserving coherence across timesteps.
We train the entire hybrid model end-to-end using classical backpropagation, optimizing the parameters $\bm{\Theta}=\{\mathbf{W}_{1},\mathbf{b}_{1},\mathbf{W}_{2},\mathbf{b}_{2}\}$ via standard optimizers, such as Adam [Kingma and Ba, 2014]. Because each $\mathbf{z}_{t}$ is real-valued, it can be used both as a per-timestep output and as a contextual embedding for soft attention in sequence-to-sequence decoding.
4 Experiments
We use the ansatz shown in Fig.1(a) (scaled to more qubits when required) as the core circuit for the QRNN. Sim et al. [2019] demonstrate experimentally that this ansatz is expressive, capable of generating strong entanglement, and able to represent a significant portion of the Hilbert space, even compared to deeper circuits built from less expressive ansätze. [^9] See Appendix B for details on the PQC design and expressibility evaluation methodology. We implement and simulate the model using TorchQuantum [Wang et al., 2022], which remains less optimized than classical toolkits due to the lack of efficient kernels for hybrid operations involving tight classical–quantum feedback, particularly in recurrent settings. Our ansatz balances expressivity, implementation simplicity, and simulation efficiency.
For $\mathbf{Measure}$ in Eq. 3, measurements are performed in each of the $\mathbf{X}, Pauli$-$\mathbf{Y}$, and Pauli-$\mathbf{Z}$ observables across all wires in the PQC, and the measurement outcomes are combined to form $\mathbf{z}_{t}$ (Eq. 3). For the feedforward network $\mathcal{F}$ (Eq. 1 and Eq. 2), we experimented with ReLU, LeakyReLU, GLU and GELU nonlinearities. [^10] GLU requires projecting to twice the output dimensionality, effectively increasing the parameter count compared to standard nonlinearities like ReLU, when all other dimensions are held constant. For both language modeling and translation, we first transform the measurement outcomes with a separate feedforward layer and use the result both for vocabulary classification and as input to the next timestep.
All experiments are run on a single A100/A30 GPU and we select the best models on the validation split across different random seeds and report the test results. The per-epoch training runtime ranges from ~4 minutes for MNIST (with 10 qubits) to ~60 minutes for language modeling (with 14 qubits). Hyperparameters shared across all the tasks include the Adam optimizer without learning rate decay $(lr=1\times 10^{-3},\ \lambda=1\times 10^{-4},\ \text{and}\ \epsilon=1\times 10^{-10})$ and dropout applied to the input at each step, with task-dependent drop rates. We apply full-sequence backpropagation without truncation, except for language modeling, where sequences are truncated to 35 tokens. No pretrained word embeddings are used. Additional hyperparameters and test set statistics (mean, min, max across runs) are provided in Appendix C. For scoRNN, we use a hidden size of 170 and the hyperparameters from Helfrich et al. [2018] are used throughout.
4.1 Sentiment Analysis
Table 1: Classification accuracy on IMDB. Qubit count $q$, total measurements $m$; or hidden state size $h$ (for RNN, LSTM and scoRNN only); embedding dimension $e$; parameter count pp. $\dagger$ indicates the LSTM in Dai and Le [2015].
[1.2pt] Model Val Test $q_{m}$ $\scriptstyle\lor$ $h$ $e$ $p$
QRNNReLU 87.25 85.37 824 100 5.2K
QRNNLeakyReLU 87.41 87.00 824 100 5.2K
QRNNGELU 87.53 86.38 824 100 5.2K
[dashed] QRNNLinear 85.37 84.21 824 100 5.2K
QRNNLinear 84.21 83.22 412 100 2.6K
RNN 87.64 86.96 50 50 5K
LSTM 88.40 86.79 25 25 5.1K
[…]
[dashed] scoRNN 84.05 83.14 170 100 31K
The IMDB sentiment dataset [Maas et al., 2011] is a balanced binary classification benchmark with 25K labeled reviews each for training and testing. The average review length is 241 tokens, with a maximum length of 2,500 tokens. We use 7.5K reviews from the training set for validation and truncate all reviews to a maximum length of 400 tokens across all models.
The hybrid model for this task follows the generic hybrid architecture described in §3.2. At the final input token, we apply an affine transformation to the measurement outcomes to produce two logits, which are used for classification via cross-entropy. Table 1 summarizes the results. QRNNLeakyReLU achieves the highest test accuracy. Ablating the classical nonlinearity (Eq. 1) degrades performance, though increasing the number of qubits in the linear model still yields some accuracy gains. Adding the nonlinearity results in a substantial improvement, outperforming all baselines. On this task, the orthogonal scoRNN underperforms other models, despite having a larger hidden state and over five times more parameters.
4.2 MNIST and Permuted-MNIST
Table 2: Classification accuracy on MNIST and pMNIST. Qubit count $q$, total measurements $m$; or hidden state size $h$ (for RNN, LSTM and scoRNN only); embedding dimension $e$; parameter count pp. $\dagger$ indicates the QRNN model of [Bausch, 2020] with 13 qubits and each digit downscaled to $4\times 4$ and binarized.
[1.2pt] MNIST pMNIST
Model Val Test Val Test $q_{m}$ $\scriptstyle\lor$ $h$ $e$ $p$
QRNNReLU 98.10 97.83 94.86 95.05 1030 28 3.9K
QRNNLeakyReLU 98.01 97.96 95.13 94.86 1030 28 3.9K
QRNNGELU 98.17 98.03 95.38 95.58 1030 28 3.9K
[dashed] QRNNLinear 97.06 96.80 94.94 94.13 1030 28 3.9K
QRNNLinear 94.31 93.87 91.10 90.55 515 28 1.3K
[…]
RNN 97.42 97.28 95.16 94.28 50 28 3.9K
LSTM 97.61 97.44 94.92 93.93 20 28 3.9K
[dashed] scoRNN 97.94 97.12 96.86 95.56 170 28 16K
We report results on the full MNIST dataset without downsampling using the same model as for IMDB, except with 10 output classes instead of binary classification. The standard pixel-by-pixel permuted MNIST (pMNIST) setup [Le et al., 2015, Arjovsky et al., 2016] requires 784 steps to process each $28\times 28$ digit, which makes simulation prohibitively slow. Here we permute the pixels of each digit first, which are then reshaped back to $28\times 28$. In both the standard and permuted cases, we use the same hyperparameters.
Table 2 shows that QRNNs with three different types of nonlinearity outperform the classical baselines on both tasks, clearly demonstrating the benefit of adding classical nonlinearities compared to the QRNNLinear models. We observe that permutation leads to an accuracy drop across all models: $2.45\%$ for QRNNGELU, $3.00\%$ for the RNN, $3.51\%$ for the LSTM, and $1.51\%$ for scoRNN, which achieves comparable performance to QRNNGELU.
4.3 Copying Memory
Figure 2: Test loss (a) and accuracy (b) for copying memory with $T=200$.
The copying memory problem tests a model’s ability to retain and recall information over long sequences [Hochreiter and Schmidhuber, 1997, Arjovsky et al., 2016]. Each input sequence has $T+20$ tokens, where the first $k=10$ are random digits from 1 to 8 ($n_{\text{classes}}$), followed by zeros, and the last 11 ($k+1$) positions are filled with the digit ‘9’ with the first ‘9’ acting as a delimiter. The model must learn to detect the delimiter and recall the original digits right after it in the output sequence. We randomly generated 5K training and 1K test samples with $T=200$ (for training efficiency of QRNNs). A random guess baseline yields a loss of $\frac{k\cdot\log(n_{\text{classes}}-1)}{T+2k}\approx 0.095,$ reflecting the expected cross-entropy when choosing uniformly from incorrect digits. On this task, QRNN-2.3K matches LSTM-168K (loss 0.07, accuracy 97%) and outperforms LSTM-2.8K (loss 0.25, accuracy 89.4%). scoRNN, specialized for this task, achieves near-perfect results, highlighting a performance gap between general-purpose and tailored models.
4.4 Word-Level Language Modeling
Table 3: PTB word-level language modeling (PPL). Qubit count $q$, total measurements $m$; or hidden state size $h$ (for RNN and LSTM only); embedding dimension $e$; parameter count pp.
[1.2pt] Model Val Test $q_{m}$ $\scriptstyle\lor$ $h$ $e$ $p$
QRNNReLU 131.81 126.69 1442 650 130K
QRNNLeakyReLU 131.41 126.58 1442 650 130K
QRNNGELU 136.62 131.07 1442 650 130K
[dashed] QRNNLeakyReLU 135.00 130.35 1030 512 78K
QRNNLeakyReLU 169.17 161.09 515 512 39K
RNN 151.96 139.13 256 256 131K
LSTM 124.22 120.30 128 128 131K
The PTB dataset [Mikolov et al., 2011] consists of 929K training tokens, 73K validation tokens, and 82K test tokens. As is standard, we use a vocabulary size of 10K, converting OOV tokens to UNK. We tested scoRNN on this task, but it did not converge to a good solution. The LSTM achieved the best result, with 120.30 perplexity (PPL), followed closely by QRNNLeakyReLU at 126.58.
4.5 Machine Translation
Table 4: Multi30K German-to-English translation (BLEU). Qubit count $q$, total measurements $m$; or hidden state size $h$ (for RNN and LSTM only); embedding dimension $e$; parameter count pp.
[1.2pt] Model Val Test $q_{m}$ $\scriptstyle\lor$ $h$ $e$ $p$
QRNNGLU 31.08 31.92 1339 512 390K
QRNNLeakyReLU 29.22 28.99 1339 512 340K
QRNNGELU 29.95 29.14 1339 512 340K
[dashed] QRNNGLU 30.16 31.51 1030 512 360K
QRNNGLU 27.63 29.66 515 512 270K
RNN 29.17 29.20 512 256 390K
LSTM 29.20 32.20 256 124 390K
Soft attentions can be implemented using various formulations, such as additive attention or dot-product attention [Luong et al., 2015], but they share the same core principle: at each decoder timestep, compute a similarity score between the current decoder state and each encoder state, normalize these scores via a softmax, and form a context vector by summation, which is then combined with the decoder’s hidden state to generate the next output token.
The attention mechanism implemented here follows the additive attention of Bahdanau et al. [2015]. At each decoding step, the decoder hidden state is concatenated with encoder outputs, passed through a tanh activation followed by a linear projection to compute alignment scores. A softmax then normalizes these scores into attention weights, with masking applied to exclude padded positions.
We applied the model to Multi30k German-to-English translation [Elliott et al., 2016], with vocabulary sizes of 19.2K for German and 10.8K for English, and an average of 11 tokens per sentence in both languages. The training set contains 29K sentence pairs, with 1K each for validation and testing.
Results in Table 4 show that QRNNGLU with 13 qubits closely matches the LSTM, followed by QRNNGLU with 10 qubits. For the QRNN, it is somewhat surprising that intermediate readouts can still support mechanisms like soft attention, since these readouts capture only partial projections of the quantum state rather than the full hidden state. This suggests that, despite intermediate measurements, sufficient information is retained and propagated across timesteps. We qualitatively interpret the learned soft alignments on a few examples where the translations required non-trivial linguistic interpretations in Appendix D.
4.6 Hidden State Gradients
Figure 3: Normalized per-timestep gradient norms $\lVert\partial\mathcal{L}/\partial\mathbf{h}_{t}\rVert_{2}$, averaged over one mini-batch containing samples of identical $T$ (batch size =16=16). Curves are normalized by the final timestep $(t=T)$ gradient to compare decay shape; higher gradient values closer to $T=0$ indicate less vanishing. (a) IMDB, $T=400$. (b) pMNIST, $T=28$.
We measure per-timestep gradient norms on IMDB ($T=400$) and pMNIST ($T=28$) by retaining gradients on the per-timestep readouts (QRNN) and hidden states (LSTM) from saved checkpoints and computing $\lVert\partial\mathcal{L}/\partial\mathbf{h}_{t}\rVert_{2}$. Gradients are averaged across samples in a mini-batch and normalized by the last-step norm $\lVert\partial\mathcal{L}/\partial\mathbf{h}_{T}\rVert_{2}$ to compare decay shape.
As shown in Fig. 3, the QRNN curves remain consistently above the LSTM on both IMDB and pMNIST, indicating less vanishing through time toward the start of the sequences. All curves start with $1.0$ at $t=T$ (normalization), but the relative elevation of the QRNN curve at earlier timesteps demonstrates more stable gradient propagation. The LSTM gradient norm decays rapidly, collapsing below $10^{-4}$ on the relatively short pMNIST sequences.
5 Discussion and Conclusion
Different quantum hardware platforms currently require distinct control stacks, and architectural choices do not translate one-to-one across devices, with factors such as native gate sets, qubit connectivity, and the implementation of mid-circuit measurements all affecting the realization of a given circuit. The aim here is not to prescribe a hardware roadmap but to analyze a hardware-realistic base case under idealized classical simulation to study the empirical properties of the architecture, where we model mid-circuit observations via projective measurements as a limiting case.
As more efficient and scalable toolchains become available (e.g., future multi-GPU toolkits based on cuQuantum [Bayraktar et al., 2023]), we anticipate more faithful simulations via ancilla-mediated schemes in which auxiliary qubits are entangled with the main circuit, measured, and reset as needed while the recurrent memory remains coherent. This aligns with mid-circuit measure-and-reset operations already supported on several platforms [DeCross et al., 2022, Lis et al., 2023, Norcia et al., 2023], although hardware implementations for large-scale sequence modeling would require fault-tolerant devices capable of sustaining long coherent recurrences and real-time classical control.
This paper bridges quantum operations and recurrent learning by introducing a new hybrid QRNN whose recurrent core is implemented as a PQC steered by a classical controller. The unitary dynamics preserve norms, promoting stable gradient propagation; the controller supplies the nonlinearity and task adaptivity needed for expressiveness; and mid-circuit measurements enable per-timestep readouts and classical feedback. As techniques improve [Abbas et al., 2023] and quantum hardware matures, the architecture provides a path toward hardware-realistic quantum models for sequential learning.
Acknowledgments
I would like to thank Bob Coecke for the research environment, Dimitri Kartsaklis, Sean Tull, and David Amaro for comments on an earlier draft.
Appendix A Quantum States and Superposition
Unlike a classical bit, a qubit exists in a superposition of the states 0 and 1 in a two-dimensional complex Hilbert space: $\lvert\psi\rangle=\alpha\lvert 0\rangle+\beta\lvert 1\rangle=\begin{bmatrix}\alpha&\beta\end{bmatrix}^{T}\in\mathbb{C}^{2}$ and $\lvert 0\rangle=\begin{bmatrix}1&0\end{bmatrix}^{T}$ and $\lvert 1\rangle=\begin{bmatrix}0&1\end{bmatrix}^{T}$ are elements of the computational basis for the Hilbert space. The coefficients $\alpha$ and $\beta$ are complex numbers referred to as the amplitudes that satisfy $|\alpha|^{2}+|\beta|^{2}=1$. For a state $|\psi\rangle=\alpha|0\rangle+\beta|1\rangle$, the probability of obtaining $|0\rangle$ is $|\alpha|^{2}$, and the probability of obtaining $|1\rangle$ is $|\beta|^{2}$.
Appendix B PQC Template
We have chosen the PQC template based on the benchmarking study in Sim et al. [2019], which evaluates 19 different parametrized quantum circuits (PQCs) up to depth 5 (i.e., the base circuit repeated up to five times and used a single PQC). Each PQC is assessed using two key metrics: expressibility and entangling capability. The architecture referred to as ansatz-14 in Sim et al. [2019] which we use here in a single layer configuration was shown to score highly on both. This gives a good balance of simulation cost and "goodness" of the PQC.
Expressibility is quantified by comparing the distribution of pairwise fidelities between states generated by the PQC to the theoretical fidelity distribution of Haar-random states, which represent uniform randomness over the composite Hilbert space (the tensor product of individual qubit spaces). Instead of generating Haar-random states directly, the method in [Sim et al., 2019] uses the analytical form of the Haar fidelity distribution as a reference. PQC output states are obtained by sampling random parameters, and their pairwise fidelities are used to construct an empirical distribution. The KL divergence between this empirical distribution and the Haar reference provides a scalar expressibility score, with lower values indicating greater expressiveness.
Appendix C Experimental Settings and Test Accuracy Statistics Across Runs
Table 5: Hyperparameters: batch size $b$, dropout rate $d$; embedding initialization $e_{init}$.
[1.2pt] Task $b$ $d$ $e_{init}$
IMDB 200 0.25 Xavier Uniform
MNIST 200 0.0  -
PTB 64 0.5 Xavier Uniform
Multi30K 64 0.25 Xavier Uniform
Table 6: Accuracy statistics on IMDB test set across 100 runs for each nonlinearity variant. Qubit count $q$, total measurements $m$; embedding dimension $e$; parameter count pp. Among all tasks, IMDB showed the greatest variability in QRNN performance across random seeds in development. This behavior may align with known sensitivities in training variational PQCs [Grant et al., 2019]. We therefore also report stats where we remove failed runs ($70\%$ accuracy, well below simple baselines such as BoW), indicated by *. For the three nonlinearities 40, 42 and 25 failed runs were observed each. The results also indicate that GELU nonlinearity reduces the sensitivity compared with the other two.
[1.2pt] Model minmin maxmax $\mu$ $min^{*}$ $\mu^{*}$ $q_{m}$ $e$ $p$
QRNNReLU 49.55 85.96 71.18 71.74 83.11 824 100 5.2K
QRNNLeakyReLU 49.63 87.00 70.23 75.77 83.44 824 100 5.2K
QRNNGELU 49.98 86.38 77.18 70.39 83.75 824 100 5.2K
While parametrized quantum circuits (PQCs) can suffer from vanishing gradients in deep or wide settings due to the barren plateau phenomenon [McClean et al., 2018], there is no general impossibility theorem that barren plateaus must occur in all parametrized quantum circuits; their presence and severity are known to depend on the ansatz, cost function, initialization, training strategy, and noise, and remain an empirical matter at practical scales. Several studies provide insights into how it arises or design principles that prevent or mitigate plateaus [Cerezo et al., 2019, Grant et al., 2019, Patti et al., 2021, Sack et al., 2022]. These results indicate that barren plateaus are not inevitable, and that careful design yields a tractable and stable training landscape in practice. In particular, some architectures such as quantum convolutional neural networks avoid barren plateaus by construction [Pesah et al., 2021], which supports the view that appropriate architectural choices can produce stable and trainable quantum models.
Table 7: Accuracy statistics on MNIST and pMNIST test sets across 50 runs for each nonlinearity variant. Qubit count $q$ and total measurements $m$; embedding dimension $e$; parameter count pp.
[1.2pt] MNIST pMNIST
Model minmin maxmax $\mu$ minmin maxmax $\mu$ $q_{m}$ $e$ $p$
QRNNReLU 97.51 98.25 97.84 94.33 95.31 94.83 1030 28 3.9K
QRNNLeakyReLU 97.42 98.15 97.88 94.33 95.38 94.80 1030 28 3.9K
QRNNGELU 97.62 98.22 97.96 94.72 95.58 95.12 1030 28 3.9K
Table 8: BLEU evaluations on the Multi30K German to English test set across 20 runs for each nonlinearity variant. Qubit count $q$, total measurements $m$; embedding dimension $e$; parameter count pp.
[1.2pt] Model minmin maxmax $\mu$ $q_{m}$ $e$ $p$
QRNNGLU 19.83 31.92 27.88 1339 512 390K
QRNNLeakyReLU 24.52 29.87 28.55 1339 512 340K
QRNNGELU 25.71 30.29 29.09 1339 512 340K
Appendix D Attention Alignments
Figure 4: Soft attention alignments produced by the QRNN encoder-decoder model.
To qualitatively analyze the model’s learned soft attention alignments we selected four sentences from test set and interpreted the hybrid model translations and alignments (Fig. 4).
We observe that the hybrid model can manage spatial and syntactic shifts while capturing clause-level structure and semantics through its measurement-driven hidden states and soft attention as well as the LSTM baseline. It is evident that the model handles compound verb constructions and semantic expansion, in sentences like “Diese Band bereitet sich auf einen Auftritt vor Publikum in einer Kirche vor” (Fig. 4(a)) and “Zwei grün gekleidete Männer bereiten in einem Restaurant Essen zu” (Fig. 4(b)), where German separable verbs—“bereitet … vor” and “bereiten … zu”—are correctly reconstructed into the English verb phrases “is preparing to perform” and “preparing”, respectively. The soft attention allowed the model to attend across non-contiguous source tokens, enabling reassembly of verb phrases. Additionally, lexical expansions such as “Publikum” $\rightarrow$ “a crowd of people” (Fig. 4(a)) and “gekleidete Männer” $\rightarrow$ “men in green outfits” (Fig. 4(b)) demonstrate contextually appropriate semantic elaboration beyond literal translation.
The model also displays syntactic reordering and clause realignment, necessitated by divergences between German and English word order. This is shown in both “Diese Band … vor Publikum … vor” and (Fig. 4(a)) “Menschen, die vor einem großen Gebäude im Kreis sitzen” (Fig. 4(c)). In the former, German’s verb-final structure is reorganized into a mid-sentence English verb phrase, while handling nested prepositional phrases. In the latter, the relative clause “die … sitzen” is compressed into the participial phrase “sitting”, dropping auxiliaries and pronouns to better fit English syntactic norms. Similarly, the location and positional phrases “im Kreis” and “vor einem großen Gebäude” are reordered into “in a circle in front of a large building”
Lastly, for multi-clause coordination, tense adaptation, and long-range dependency tracking, as seen in “Acht Männer spielen auf der Bühne, während ein Gitarrist im Scheinwerferlicht spielt” (Fig. 4(d)). The model successfully disentangles two coordinated clauses and renders them with the correct English conjunction “while”, while adjusting verb forms from German’s uniform “spielen” to “are playing” and “plays”, based on subject plurality. Finally, this ability to flexibly adapt clause boundaries and maintain coherence is also reflected in the “Menschen … im Kreis sitzen” example (Fig. 4(c)), where the model tracks relative clause dependencies and maps them onto compact English constructions.

### v2

@@
 Abstract
-We present a hybrid quantum-classical recurrent neural network (QRNN) architecture in which the entire recurrent core is realized as a parametrized quantum circuit (PQC) controlled by a classical feedforward network.
-The hidden state is the quantum state of an $n$-qubit PQC, residing in an exponentially large Hilbert space $\mathbb{C}^{2^{n}}$.
+We present a hybrid quantum-classical recurrent neural network (QRNN) architecture in which the recurrent core is realized as a parametrized quantum circuit (PQC) controlled by a classical feedforward network.
+The hidden state is the quantum state of an $n$-qubit PQC in an exponentially large Hilbert space $\mathbb{C}^{2^{n}}$, which serves as a coherent recurrent quantum memory.
 The PQC is unitary by construction, making the hidden-state evolution norm-preserving without external constraints.
-At each timestep, mid-circuit readouts are combined with the input embedding and processed by the feedforward network, which provides explicit classical nonlinearity.
+At each timestep, mid-circuit Pauli expectation-value readouts are combined with the input embedding and processed by the feedforward network, which provides explicit classical nonlinearity.
 The outputs parametrize the PQC, which updates the hidden state via unitary dynamics.
-The QRNN is compact and physically consistent, and it unifies (i) unitary recurrence as a high-capacity memory, (ii) partial observation via mid-circuit measurements, and (iii) nonlinear classical control for input-conditioned parametrization.
-We evaluate the model in simulation with up to 14 qubits on sentiment analysis, MNIST, permuted MNIST, copying memory, and language modeling, adopting projective measurements as a limiting case to obtain mid-circuit readouts while maintaining a coherent recurrent quantum memory.
-We further devise a soft attention mechanism over the mid-circuit readouts in a sequence-to-sequence model and show its effectiveness for machine translation.
+The QRNN is compact and physically consistent, and it unifies (i) unitary recurrence as a high-capacity memory, (ii) partial observation via mid-circuit readouts, and (iii) nonlinear classical control for input-conditioned parametrization.
+We evaluate the model in simulation with up to 14 qubits on sentiment analysis, MNIST, permuted MNIST, copying memory, and language modeling.
+For sequence-to-sequence learning, we further devise a soft attention mechanism over the mid-circuit readouts and show its effectiveness for machine translation.
 To our knowledge, this is the first model (RNN or otherwise) grounded in quantum operations to achieve competitive performance against strong classical baselines across a broad class of sequence-learning tasks.
@@
 This means that more complex sequences may exceed what the hidden state can encode, forcing the model to compress or forget.
-The introduction of the Transformer [Vaswani et al., 2017] appeared to obviate explicit recurrence by bypassing the hidden-state bottleneck.
-However, recent work shows that recurrent inductive bias remains highly competitive and provides representational advantages not matched by Transformers [Gu and Dao, 2023, Orvieto et al., 2023, Bhattamishra et al., 2024, Beck et al., 2024].
 Another challenge in training RNNs is the vanishing and exploding gradient problem [Bengio et al., 1994, Hochreiter and Schmidhuber, 1997], which arises from repeated multiplication through the recurrent Jacobian.
@@
 These models perform well on synthetic tasks, but their results on broader benchmarks vary.
+The introduction of the Transformer [Vaswani et al., 2017] appeared to obviate explicit recurrence by bypassing the hidden-state bottleneck.
+However, recent work shows that recurrent inductive bias remains highly competitive and provides representational advantages not matched by Transformers [Gu and Dao, 2023, Orvieto et al., 2023, Bhattamishra et al., 2024, Beck et al., 2024].
 With the advancement of quantum computing [Arute et al., 2019, Acharya et al., 2024, Reichardt et al., 2024, DeCross et al., 2025], parametrized quantum circuits (PQCs), which are a core component of variational hybrid quantum-classical models, have concurrently emerged as an alternative mechanism for function approximation [Benedetti et al., 2019, Du et al., 2019, Bondesan and Welling, 2020, Pérez-Salinas et al., 2021, Schuld et al., 2021, Yu et al., 2024b].
@@
 The feedforward network $\mathcal{F}$ takes as input the concatenation of the mid-circuit readout vector from the previous timestep and the current input $(\mathbf{z}_{t-1}\mathpunct{:}\mathbf{x}_{t})$, and outputs $\bm{\theta}_{t}\in\mathbb{R}^{16}$ containing the 16 rotation angles that parametrize the PQC shown in (a) as $U(\bm{\theta}_{t})$, which acts on the quantum state propagated from the previous step.
-Refer to caption denotes qubit measurements.
-In this work, we present a hybrid quantum–classical recurrent neural network (QRNN) architecture grounded in quantum operations, in which the entire recurrent core is realized as a PQC.
-The hidden state is the quantum state of the PQC, residing in an exponentially large Hilbert space.
-A classical feedforward network parametrizes and steers the quantum computation, introducing nonlinearity through mid-circuit readouts, which are realized in simulation as a limiting case via projective measurements.
-This avoids emulating nonlinearity through linear quantum dynamics, leaving the PQC strictly for coherent unitary evolution.
-Fig. 2 illustrates both the PQC (with four qubits shown for illustration) and the unrolled QRNN:
-  • At each timestep $t$, the input is mapped to a classical embedding $\mathbf{x}_{t}$ via a learnable embedding layer.
-  • A classical feedforward network $\mathcal{F}$ takes as input the concatenation of the readout vector $\mathbf{z}_{t-1}$ (outputs from all measurements at timestep t1t-1) and the current input $\mathbf{x}_{t}$.
+Refer to caption denotes qubit expectation-value readouts.
+In this work, we present a hybrid quantum-classical QRNN in which the recurrent core is realized as a PQC controlled by a classical feedforward network.
+The hidden state is the quantum state of the $n$-qubit PQC, residing in an exponentially large Hilbert space $\mathbb{C}^{2^{n}}$, which provides a high-capacity coherent quantum memory.
+Nonlinearity is introduced through classical computation rather than approximated within the quantum circuit, leaving the PQC strictly for unitary evolution of the hidden state.
+Fig.
+2 illustrates both the PQC (with four qubits shown for illustration) and the unrolled QRNN:
+  • At timestep $t$, the input is mapped to a learnable representation $\mathbf{x}_{t}$ via an embedding layer.
+  • A classical feedforward network $\mathcal{F}$ takes as input the concatenation of $\mathbf{z}_{t-1}$ (readout outputs at timestep t1t-1) and the current input $\mathbf{x}_{t}$.
 It outputs the PQC parameters $\bm{\theta}_{t}$ which configure the PQC with a fixed gate layout (Fig.
@@
 Residing in an exponentially large Hilbert space, this state persists across timesteps and provides the model’s core recurrent memory.
-  • The mid-circuit readout $\mathbf{z}_{t}$ (or the final readout at the end of the sequence) is a classical feature vector obtained from the quantum state via measurements and is used: (i) as recurrent feedback $\mathbf{z}_{t-1}$ at timestep $t$, and (ii) as the input to task-specific classical layers.
+  • The mid-circuit readout $\mathbf{z}_{t}$ (or the final readout at the end of the sequence) is a real-valued vector obtained from the quantum state via Pauli expectation-value readouts and is used: (i) as recurrent feedback $\mathbf{z}_{t-1}$ at timestep $t$, and (ii) as the input to task-specific classical layers.
 We develop the models on GPUs, allowing us to simulate and train quantum recurrence via classical backpropagation, with the expectation that such models will become classically unsimulatable as the number of qubits increases.
-To our knowledge, this is the first model grounded in quantum operations (RNN or otherwise) demonstrated in classical simulation with up to 14 qubits across six realistic sequence-modeling tasks, achieving competitive performance with LSTM and the scaled Cayley orthogonal scoRNN designed for norm preservation [Helfrich et al., 2018].
-Experiments also show that classical nonlinear control and feedback are effective, with the nonlinear variants outperforming their linear counterparts, and that the unitary quantum recurrent core maintains more stable gradients than LSTMs (§4.6).
+To our knowledge, this is the first model (RNN or otherwise) grounded in quantum operations demonstrated in classical simulation with up to 14 qubits across six realistic sequence-modeling tasks, achieving competitive performance with LSTM and the scaled Cayley orthogonal scoRNN designed for norm preservation [Helfrich et al., 2018].
+Experiments also show that classical nonlinear control and feedback are effective, with nonlinear variants outperforming their linear counterparts, and that the unitary quantum recurrent core maintains more stable gradients than LSTMs on sequences of up to 400 tokens (§4.6).
 The QRNN is motivated in part by the memory and gradient problems of RNNs, but its main aim is to explore a hybrid quantum–classical recurrent model in an idealized proof of principle that allows us to study its computational behavior under best-case conditions across a broad class of sequence learning tasks.
@@
 2 Related Work
-Bausch [Bausch, 2020] developed a QRNN with a persistent quantum memory based on quantum neurons [Cao et al., 2017].
-However, nonlinearities are emulated within PQCs using probabilistic repeat-until-success circuit execution and postselection steps.
-The inherently linear nature of quantum computation constrains such emulation, and the available forms of nonlinearity remain limited [Yan et al., 2020, Moreira et al., 2023, Zi et al., 2024].
+Bausch [2020] proposes a QRNN whose recurrence is implemented by iterating a quantum cell built from “quantum neurons”.
+Nonlinearity is induced inside the circuit via measurement-and-postselection primitives (repeat-until-success), together with amplitude amplification to make the procedure near-deterministic.
+Consequently, the per-timestep update is not strictly unitary and introduces measurement back-action on the memory.
+Because the dynamics remain linear between measurements, the effective nonlinear maps achievable by such measurement-based schemes are structurally constrained, and the repertoire of admissible nonlinearities remains limited [Yan et al., 2020, Moreira et al., 2023, Zi et al., 2024].
 The so-called QLSTMs embed PQCs into the gating mechanisms of classical LSTMs [Chen et al., 2020, Yu et al., 2024a, Ubale et al., 2025], replacing dense layers in the LSTM gates with PQCs.
@@
 [2023] and Siemaszko et al.
-[2023] also model recurrences with PQCs, while supporting per-timestep readouts, but they rely entirely on linear quantum dynamics of the PQC without explicit nonlinearities or classical control.
+[2023] also model recurrences with PQCs and support per-timestep readouts, but they rely entirely on linear quantum dynamics without other nonlinearity or classical control.
 Experiments with the existing models have focused on domain-specific tasks such as fraud detection [Ubale et al., 2025], low-resource text classification [Yu et al., 2024a], or scaled-down MNIST [Bausch, 2020, Siemaszko et al., 2023].
@@
 1(a) consists of entirely parametrized gates.
-Measurements.
-To probe the quantum state of a PQC, we can perform measurements to obtain real-valued readouts.
-These readouts provide partial observations of the state, and any required number of measurements, on any of the qubits, can be combined for downstream tasks.
-For instance, a measurement through the Pauli-$\mathbf{Z}$ observable with the unitary
-$\mathbf{Z}=\begin{pmatrix}1&0\\ 0&-1\end{pmatrix}$
-assigns scalar values (e.g., +1+1 for $|0\rangle$ and 1-1 for $|1\rangle$) in the computational basis in the single-qubit case.
-For a general quantum state however, the outcome is probabilistic: it yields +1+1 with probability $|\alpha|^{2}$ and 1-1 with probability $|\beta|^{2}$.
-The expectation value of this measurement is given by $|\alpha|^{2}-|\beta|^{2}$, which can be used as a real-valued readout in hybrid quantum-classical models.
-Although the readouts obtained via measurement are a nonlinear function of the gate parameters, particularly those used in parametrized rotation gates such as $\mathbf{RX}$, the resulting nonlinearity is generally weak (§4).
+Readouts via Pauli expectations.
+A succinct way to summarize a quantum state is via expectation values of Pauli observables.
+For each qubit $k$, the Pauli operators $X_{k},Y_{k},Z_{k}$ are the three fundamental single-qubit observables that act on that qubit (with identity on all other qubits).
+For example, the Pauli-$Z$ operator measures the computational basis with
+$Z\;=\;\begin{pmatrix}1&0\\[2.0pt] 0&-1\end{pmatrix}$,
+while $X$ and $Y$ measure superposition states in orthogonal bases.
+For a single qubit $\lvert\psi\rangle=\alpha\lvert 0\rangle+\beta\lvert 1\rangle$, the Pauli-$Z$ expectation is $\langle Z\rangle=\langle\psi\lvert Z\rvert\psi\rangle=|\alpha|^{2}-|\beta|^{2}\in[-1,1]$.
+For an $n$-qubit state, per-qubit expectations $\{\langle X_{k}\rangle,\langle Y_{k}\rangle,\langle Z_{k}\rangle\}_{k=1}^{n}$ and correlation terms (Pauli strings) $\langle P_{1}\otimes\cdots\otimes P_{n}\rangle$ with $P_{\ell}\in\{I,X,Y,Z\}$ yield bounded, hardware-agnostic real-valued summaries useful for analysis and downstream modeling.
+Although expectation-value readouts are nonlinear functions of parametrized rotation angles (e.g., in $\mathbf{RX}$ gates), the resulting nonlinearity is generally insufficient on its own (§4).
 3.2 Hybrid Model
@@
 1(a), which is controlled by a classical feedforward network and evolved by applying the unitary gates.
-Let $\mathbf{x}_{t}$ be the input embedding at timestep $t$, and let $\mathbf{z}_{t-1}$ be the measurement-based readout from the previous timestep.
-In the most generic form of the hybrid model, [^7] We may add extra transformations to the measurement outcomes before classifications or feeding them to the next step; see §4.
+Let $\mathbf{x}_{t}$ be the input embedding at timestep $t$, and let $\mathbf{z}_{t-1}$ be the readout vector from the previous timestep.
+In the most generic form of the hybrid model, [^7] Extra transformations may be applied to the readouts before classifications or feeding them to the next step (for some tasks); see §4.
 the two are concatenated into a single vector $\mathbf{u}_{t}=(\mathbf{z}_{t-1}\mathpunct{:}\mathbf{x}_{t})$ and passed through a classical feedforward network $\mathcal{F}$ with one hidden layer and a nonlinearity.
@@
 1(a)).
-The PQC itself is defined by a unitary operator $U(\bm{\theta}_{t})$, parametrized by $\bm{\theta}_{t}$.
-[^8] We use $U(\bm{\theta}_{t})$ to denote all unitary operations composed of multiple parametrized gates, each acting on one or more qubits with parameters drawn from $\bm{\theta}_{t}$.
-Applying the gates in $U(\bm{\theta}_{t})$ to the quantum state $\mathbf{h}_{t-1}=|\psi_{t-1}\rangle$ yields the updated state $\mathbf{h}_{t}=U(\bm{\theta}_{t})\,|\psi_{t-1}\rangle$.
-The updated state is then measured to obtain a classical readout vector
-$\mathbf{z}_{t}=\textbf{Measure}(\mathbf{h}_{t}),$ (3)
-which serves as a proxy for the quantum state and is combined with the next input $\mathbf{x}_{t+1}$ to evolve the recurrence.
-Mid-circuit measurements are simulated without collapsing the quantum state via projective measurements, thereby preserving coherence across timesteps.
+The PQC itself is defined by a unitary operator $U(\bm{\theta}_{t})$ parametrized by $\bm{\theta}_{t}$.
+[^8] Here $U(\bm{\theta}_{t})$ denotes the product of the circuit’s parametrized gates, each acting on one or more qubits with its parameter drawn from $\bm{\theta}_{t}$.
+Applying $U(\bm{\theta}_{t})$ to the prior state $\mathbf{h}_{t-1}=\lvert\psi_{t-1}\rangle$ yields the updated state
+$\mathbf{h}_{t}=U(\bm{\theta}_{t})\,\lvert\psi_{t-1}\rangle.$
+A classical readout vector is then computed from Pauli expectation values:
+$\mathbf{z}_{t}=\mathrm{Readout}(\mathbf{h}_{t})=\big(\,\langle X_{1}\rangle_{t},\ \langle Y_{1}\rangle_{t},\ \langle Z_{1}\rangle_{t},\ \ldots,\ \langle X_{n}\rangle_{t},\ \langle Y_{n}\rangle_{t},\ \langle Z_{n}\rangle_{t}\,\big),$ (3)
+where $\langle P_{k}\rangle_{t}=\langle\psi_{t}\lvert P_{k}\rvert\psi_{t}\rangle$ for $P_{k}\in\{X_{k},Y_{k},Z_{k}\}$ and $\lvert\psi_{t}\rangle=\mathbf{h}_{t}$.
+[^9] For computational efficiency, we use only single-qubit expectations rather than multi-qubit Pauli strings.
+Serving as a proxy for the quantum state, $\mathbf{z}_{t}$ is concatenated with the next input $\mathbf{x}_{t+1}$ to produce $\bm{\theta}_{t+1}$ via the classical controller.
+Because the Pauli expectation-value readouts do not alter the state, coherence of the quantum state is preserved across timesteps, allowing it to function as a coherent recurrent memory.
 We train the entire hybrid model end-to-end using classical backpropagation, optimizing the parameters $\bm{\Theta}=\{\mathbf{W}_{1},\mathbf{b}_{1},\mathbf{W}_{2},\mathbf{b}_{2}\}$ via standard optimizers, such as Adam [Kingma and Ba, 2014].
-Because each $\mathbf{z}_{t}$ is real-valued, it can be used both as a per-timestep output and as a contextual embedding for soft attention in sequence-to-sequence decoding.
+For sequence-to-sequence learning, $\mathbf{z}_{t}$ provides per-timestep outputs and serves as contextual embeddings for soft attention decoding.
 4 Experiments
@@
 Our ansatz balances expressivity, implementation simplicity, and simulation efficiency.
-For $\mathbf{Measure}$ in Eq.
-3, measurements are performed in each of the $\mathbf{X}, Pauli$-$\mathbf{Y}$, and Pauli-$\mathbf{Z}$ observables across all wires in the PQC, and the measurement outcomes are combined to form $\mathbf{z}_{t}$ (Eq.
-3).
 For the feedforward network $\mathcal{F}$ (Eq.
@@
 [^11] GLU requires projecting to twice the output dimensionality, effectively increasing the parameter count compared to standard nonlinearities like ReLU, when all other dimensions are held constant.
-For both language modeling and translation, we first transform the measurement outcomes with a separate feedforward layer and use the result both for vocabulary classification and as input to the next timestep.
+For both language modeling and translation, we first transform the readouts with a separate feedforward layer and use the result both for vocabulary classification and as input to the next timestep.
 All experiments are run on a single A100/A30 GPU and we select the best models on the validation split across different random seeds and report the test results.
@@
 Table 1: Classification accuracy on IMDB.
-Qubit count $q$, total measurements $m$; or hidden state size $h$ (for RNN, LSTM and scoRNN only); embedding dimension $e$; parameter count pp.
+Qubit count $q$, total readouts $m$; or hidden state size $h$ (for RNN, LSTM and scoRNN only); embedding dimension $e$; parameter count pp.
 $\dagger$ indicates the LSTM in Dai and Le [2015].
@@
 The hybrid model for this task follows the generic hybrid architecture described in §3.2.
-At the final input token, we apply an affine transformation to the measurement outcomes to produce two logits, which are used for classification via cross-entropy.
+At the final input token, we apply an affine transformation to the readouts to produce two logits, which are used for classification via cross-entropy.
 Table 1 summarizes the results.
@@
 On this task, the orthogonal scoRNN underperforms other models, despite having a larger hidden state and over five times more parameters.
-4.2 MNIST and Permuted-MNIST
+4.2 MNIST and Permuted MNIST
 Table 2: Classification accuracy on MNIST and pMNIST.
-Qubit count $q$, total measurements $m$; or hidden state size $h$ (for RNN, LSTM and scoRNN only); embedding dimension $e$; parameter count pp.
+Qubit count $q$, total readouts $m$; or hidden state size $h$ (for RNN, LSTM and scoRNN only); embedding dimension $e$; parameter count pp.
 $\dagger$ indicates the QRNN model of [Bausch, 2020] with 13 qubits and each digit downscaled to $4\times 4$ and binarized.
@@
 4.4 Word-Level Language Modeling
-Table 3: PTB word-level language modeling (PPL). Qubit count $q$, total measurements $m$; or hidden state size $h$ (for RNN and LSTM only); embedding dimension $e$; parameter count pp.
+Table 3: PTB word-level language modeling (PPL). Qubit count $q$, total readouts $m$; or hidden state size $h$ (for RNN and LSTM only); embedding dimension $e$; parameter count pp.
 [1.2pt] Model Val Test $q_{m}$ $\scriptstyle\lor$ $h$ $e$ $p$
@@
 4.5 Machine Translation
-Table 4: Multi30K German-to-English translation (BLEU). Qubit count $q$, total measurements $m$; or hidden state size $h$ (for RNN and LSTM only); embedding dimension $e$; parameter count pp.
+Table 4: Multi30K German-to-English translation (BLEU). Qubit count $q$, total readouts $m$; or hidden state size $h$ (for RNN and LSTM only); embedding dimension $e$; parameter count pp.
 [1.2pt] Model Val Test $q_{m}$ $\scriptstyle\lor$ $h$ $e$ $p$
@@
 For the QRNN, it is somewhat surprising that intermediate readouts can still support mechanisms like soft attention, since these readouts capture only partial projections of the quantum state rather than the full hidden state.
-This suggests that, despite intermediate measurements, sufficient information is retained and propagated across timesteps.
+This suggests that, despite mid-circuit readouts, sufficient information is retained and propagated across timesteps.
 We qualitatively interpret the learned soft alignments on a few examples where the translations required non-trivial linguistic interpretations in Appendix D.
@@
 5 Discussion and Conclusion
-Different quantum hardware platforms currently require distinct control stacks, and architectural choices do not translate one-to-one across devices, with factors such as native gate sets, qubit connectivity, and the implementation of mid-circuit measurements all affecting the realization of a given circuit.
-The aim here is not to prescribe a hardware roadmap but to analyze a hardware-realistic base case under idealized classical simulation to study the empirical properties of the architecture, where we model mid-circuit observations via projective measurements as a limiting case.
+Different quantum hardware platforms currently require distinct control stacks, and architectural choices do not translate one-to-one across devices, with factors such as native gate sets, qubit connectivity, and the implementation of mid-circuit readout all affecting the realization of a given circuit.
+The aim here is not to prescribe a hardware roadmap but to analyze a hardware-realistic base case under idealized classical simulation to study the empirical properties of the architecture, where we model mid-circuit observations via expectation-value readouts.
 As more efficient and scalable toolchains become available (e.g., future multi-GPU toolkits based on cuQuantum [Bayraktar et al., 2023]), we anticipate more faithful simulations via ancilla-mediated schemes in which auxiliary qubits are entangled with the main circuit, measured, and reset as needed while the recurrent memory remains coherent.
@@
 This paper bridges quantum operations and recurrent learning by introducing a new hybrid QRNN whose recurrent core is implemented as a PQC steered by a classical controller.
-The unitary dynamics preserve norms, promoting stable gradient propagation; the controller supplies the nonlinearity and task adaptivity needed for expressiveness; and mid-circuit measurements enable per-timestep readouts and classical feedback.
+The unitary dynamics preserve norms, promoting stable gradient propagation; mid-circuit, per-timestep readouts inject task adaptability; and the classical controller supplies the nonlinearity and feedback for expressiveness.
 As techniques improve [Abbas et al., 2023] and quantum hardware matures, the architecture provides a path toward hardware-realistic quantum models for sequential learning.
@@
 Table 6: Accuracy statistics on IMDB test set across 100 runs for each nonlinearity variant.
-Qubit count $q$, total measurements $m$; embedding dimension $e$; parameter count pp.
+Qubit count $q$, total readouts $m$; embedding dimension $e$; parameter count pp.
 Among all tasks, IMDB showed the greatest variability in QRNN performance across random seeds in development.
@@
 In particular, some architectures such as quantum convolutional neural networks avoid barren plateaus by construction [Pesah et al., 2021], which supports the view that appropriate architectural choices can produce stable and trainable quantum models.
-Table 7: Accuracy statistics on MNIST and pMNIST test sets across 50 runs for each nonlinearity variant. Qubit count $q$ and total measurements $m$; embedding dimension $e$; parameter count pp.
+Table 7: Accuracy statistics on MNIST and pMNIST test sets across 50 runs for each nonlinearity variant. Qubit count $q$ and total readouts $m$; embedding dimension $e$; parameter count pp.
 [1.2pt] MNIST pMNIST
@@
 QRNNGELU 97.62 98.22 97.96 94.72 95.58 95.12 1030 28 3.9K
-Table 8: BLEU evaluations on the Multi30K German to English test set across 20 runs for each nonlinearity variant. Qubit count $q$, total measurements $m$; embedding dimension $e$; parameter count pp.
+Table 8: BLEU evaluations on the Multi30K German to English test set across 20 runs for each nonlinearity variant. Qubit count $q$, total readouts $m$; embedding dimension $e$; parameter count pp.
 [1.2pt] Model minmin maxmax $\mu$ $q_{m}$ $e$ $p$
@@
 To qualitatively analyze the model’s learned soft attention alignments we selected four sentences from test set and interpreted the hybrid model translations and alignments (Fig. 4).
-We observe that the hybrid model can manage spatial and syntactic shifts while capturing clause-level structure and semantics through its measurement-driven hidden states and soft attention as well as the LSTM baseline.
+We observe that the hybrid model can manage spatial and syntactic shifts while capturing clause-level structure and semantics through its readout-based hidden states and soft attention as well as the LSTM baseline.
 It is evident that the model handles compound verb constructions and semantic expansion, in sentences like “Diese Band bereitet sich auf einen Auftritt vor Publikum in einer Kirche vor” (Fig.